OpenStudy (anonymous):

what are the solutions for x in the proportion 5x-6/x=2x/2

hero (hero):

is it: \[\frac{5x - 6}{x} = \frac{2x}{2}\]

OpenStudy (anonymous):

yes

hero (hero):

Okay, so cross multiply. Let me know what you get afterwards.

OpenStudy (anonymous):

10x-12=2x^2

hero (hero):

Okay, good so you have \(10x - 12 = 2x^2\) Now, create a quadratic with \(2x^2\) as the leading term by subtracting 10x from both sides, then adding 12 to both sides. .

OpenStudy (anonymous):

i still dont understand

hero (hero):



OpenStudy (anonymous):

so the answer is (2,3)

hero (hero):

Good job. May I ask, what didn't you understand about my steps beforehand?